Crusty’s pizza company sells its pepperoni pizza for $18.75. It costs Crusty’s $9.25 to make each pepperoni pizza. Crusty’s also has fixed costs of $3500 per month. How many pepperoni pizzas must Crusty’s sell each month to break even (round to the nearest full pizza)?

A.

368

B.

187

C.

378

D.

215
Answer #1

A. 368

Let the number of pizzas to be sold each month to break even be X, then:

3500 + 9.25*X = 18.75*X

9.5*X = 3500

X = 368 Approximately